OpenStudy (anonymous):

5 over 1+i write the expression as a complex number in standard form

OpenStudy (cggurumanjunath):

rationalise the denominator !

OpenStudy (kc_kennylau):

\[\frac5{1+i}=\frac{5(1-i)}{(1+i)(1-i)}=\frac{5-i}2=2.5-0.5i\]

OpenStudy (pixiedust1):

WELCOME TO OPEN STUDY :D